# Learn by Reasoning: Analogical Weight Generation for Few-Shot Class-Incremental Learning

### v1

@@ Abstract
Few-shot class-incremental Learning (FSCIL) enables models to learn new classes from limited data while retaining performance on previously learned classes. Traditional FSCIL methods often require fine-tuning parameters with limited new class data and suffer from a separation between learning new classes and utilizing old knowledge. Inspired by the analogical learning mechanisms of the human brain, we propose a novel analogical generative method. Our approach includes the Brain-Inspired Analogical Generator (BiAG), which derives new class weights from existing classes without parameter fine-tuning during incremental stages. BiAG consists of three components: Weight Self-Attention Module (WSA), Weight &\&& Prototype Analogical Attention Module (WPAA), and Semantic Conversion Module (SCM). SCM uses Neural Collapse theory for semantic conversion, WSA supplements new class weights, and WPAA computes analogies to generate new class weights. Experiments on miniImageNet, CUB-200, and CIFAR-100 datasets demonstrate that our method achieves higher final and average accuracy compared to SOTA methods.
Index Terms:
Few-shot class-incremental learning, Analogical learning, Brain-inspired model, Classifier weight generation, Attention mechanism
Figure 1: An example of human analogical learning mechanism and the proposed Brain-Inspired Analogical Generator (BiAG). (a) Humans can efficiently form concepts of new categories through analogical reasoning. For example, when encountering a new animal like a panda, one may relate its visual characteristics to previously known categories—such as the body shape of a bear and the black-and-white color pattern of a zebra. This analogy enables the construction of a concept of ”panda” even from limited samples. (b) Inspired by this cognitive mechanism, our BiAG module analogically generates classification weights for new classes by leveraging prototypes and weights from old classes along with a learnable query derived from new class knowledge. This allows efficient integration of new categories in few-shot incremental learning.
@@ I Introduction
Conventional deep learning models have achieved remarkable success in various computer vision and pattern recognition tasks. However, these models are typically developed under a closed-world assumption, where all categories are predefined, and ample labeled data is available for each class. In practice, this assumption rarely holds. New classes may appear continuously over time, while labeled data is often expensive or difficult to obtain. To address this limitation, Class-Incremental Learning (CIL) has been proposed to enable models to learn new categories sequentially while retaining prior knowledge. Although effective in fully supervised settings, most existing CIL methods assume sufficient data for each new class and often require storing exemplars from past sessions. These assumptions greatly limit their applicability in dynamic and data-constrained environments.
@@
In many real-world scenarios, such as identifying rare species, adapting to evolving user behavior, or updating clinical diagnosis systems, new categories often emerge under tight time constraints and with very limited data. In such settings, retraining models from scratch or storing all historical data is often infeasible. Moreover, traditional CIL methods tend to overfit to the few new samples or suffer from forgetting previously learned knowledge. To address these challenges, Few-Shot Class-Incremental Learning (FSCIL) has been proposed [1]. FSCIL requires models to learn new classes from only a handful of labeled examples while maintaining performance on old classes. Typically, the model is trained in a base session with sufficient data, followed by multiple incremental sessions with only a few examples per new class. Data from previous sessions becomes unavailable, and the model must perform classification over all seen classes. This setting emphasizes the need for methods that can effectively balance stability and plasticity under strict data and time constraints.
FSCIL faces two major challenges: overfitting caused by the limited training samples of new classes and catastrophic forgetting of old classes [1]. Existing methods can be broadly categorized into unified learning approaches and sequential learning approaches. Unified learning methods aim to jointly optimize for both new and old classes by leveraging retained old knowledge. For example, meta-learning-based methods [2, 3] enhance adaptability by utilizing knowledge from related tasks, while replay-based methods [4] mitigate forgetting by replaying previously learned information. Dynamic network structures [5] employ adaptive graph models to better integrate new and old knowledge. On the other hand, sequential learning approaches treat the learning of new classes and the retention of old knowledge as separate tasks. Some of these methods focus on efficient representation learning to address the challenge of limited data [6][7][8]. Despite their merits, both approaches suffer from common limitations. First, they often rely on fine-tuning during incremental stages, which increases the risk of overfitting to the limited samples of new classes. Second, these methods treat the learning of new classes and the retention of old knowledge as separate tasks, resulting in a disconnect between the utilization of new and old knowledge. This separation makes it challenging to achieve a balanced integration of old knowledge retention and new information acquisition.
@@
Inspired by the analogical learning mechanism of the human brain, we propose a novel analogical generation method that derives new class weights from existing ones. Our method sufficiently utilizes prior knowledge during incremental sessions while eliminating the need for parameter training. The framework includes a feature extractor and a Brain-Inspired Analogical Generator (BiAG). The BiAG takes the prototypes (the centric or mean feature of an identical class) of new and old classes, as well as the classification weights of old classes, as inputs to the module, and outputs the generated classification weights of the new classes. The BiAG comprises three core components: the Semantic Conversion Module (SCM), the Weight Self-Attention Module (WSA), and the Weight & Prototype Analogical Attention Module (WPAA). The SCM facilitates smooth transitions between prototypes and classification weights based on Neural Collapse theory. The WSA enhances old knowledge utilization by supplementing new class weights, while the WPAA draws analogies between old and new classes, reorganizing old weights to generate accurate new class weights. Building upon this foundation, our framework is structured into two distinct phases: the base session training and incremental session inference phases. During the base session, the feature extractor is trained on the base classes to obtain initial classification weights and prototypes, thereby constructing the knowledge base for analogical learning. Subsequently, the BiAG is trained to cultivate the ability to generate weights for novel classes through simulated incremental learning, wherein we randomly divide the base classes into old and new classes and train the model to generate weights for new classes. In the incremental inference phase, as illustrated in Fig. 1(b), the BiAG generates new class weights by leveraging the prototypes of the new classes alongside the prototypes and weights of the old classes.
@@ II-A Class-Incremental Learning (CIL)
Class-incremental learning (CIL) enables models to continually recognize new classes while preserving knowledge of previously learned ones. The main challenge lies in catastrophic forgetting, where learning new tasks interferes with old knowledge. To address this, a variety of strategies have been developed. Regularization-based methods [14][15][16] penalize changes to important parameters to preserve past knowledge. For example, EWC [16] leverages Fisher Information to identify critical weights. More recent efforts introduce forgetting-aware directions and equilibrium constraints to improve stability [17]. Replay-based methods [18][19][20][21] retain exemplars or use generative models [22] to rehearse previous classes. To overcome memory and privacy concerns, recent approaches replace exemplars with mixed features and auxiliary classes [23] or use semantic mapping and background calibration to improve alignment [24]. Statistical sampling has also been proposed to simulate past data distributions without real samples [25]. Parameter isolation strategies [26][27][28][29]assign different parameter subsets to different tasks, avoiding interference. PackNet [27], for instance, uses weight pruning and reallocation to accommodate new tasks efficiently. Dynamic architecture methods [30][31][32][33] incrementally grow the network to handle new tasks, though often at the cost of increased complexity. Recent improvements include lightweight incremental frameworks [34] and memory-boosted transformers [35]. In addition, several works enhance representation quality to improve learning efficiency. For example, feature expansion and robust training techniques have proven effective in exemplar-free settings [36], while others leverage external semantic knowledge to guide feature learning and improve generalization [37].
@@ II-B Few-Shot Class-Incremental Learning (FSCIL)
Few-shot class-incremental learning (FSCIL) aims to learn new classes from few samples over time without forgetting old ones. Early works such as TOPIC [1] and CEC [5] adopt feature structure preservation or graph-based adaptation. Meta-learning-based methods like MetaFSCIL [3] and Self-promoted [38] improve adaptability via task sampling and prototype refinement. SPPR [39] and IDLVQ [40] enhance prototype-based learning through episodic sampling and vector quantization. To address representation degradation, FACT [41] pre-allocates virtual prototypes, while Data-free Replay [7] replays knowledge to balance class learning. LIMIT [42] and CLOM [43] employ transformers and margin constraints to align semantics and reduce overfitting. GKEAL [44] uses Gaussian kernels for analytic learning, and BiDist [45] distills knowledge from dual teachers. SVAC [8] enhances generalization via virtual classes. Dynamic model-based approaches include SubNet [46], WaRP [47], and KRRM [48], each tailoring model structures for incremental adaptation. NC-FSCIL [49] pre-assigns classifiers based on Neural Collapse, while ALICE [50] and OrCo [51] improve intra-class compactness and generalization via angular loss or orthogonality. YourSelf [52] uses pseudo-label self-training without fine-tuning. While ALFSCIL [53] explores analogical learning, it only leverages prototypes and still relies on fine-tuning. In contrast, our method, BiAG, fully utilizes both prototypes and old class weights to analogically generate new classifiers in a scalable and tuning-free manner.
@@ II-C Analogical Learning Mechanism
Analogical learning is a fundamental aspect of human cognition, allowing individuals to transfer knowledge from familiar domains (source) to novel situations (target) by recognizing relational similarities. This mechanism supports efficient and flexible learning and has been identified by cognitive science as a key driver of human intelligence [9, 10, 54]. According to Gentner’s Structure-Mapping Theory [55], analogical reasoning involves retrieving relevant experiences, aligning relational structures (mapping), projecting inferences, evaluating consistency, and re-representing knowledge when necessary.
Neurologically, analogical reasoning is supported by a distributed network involving the hippocampus, prefrontal cortex (PFC), and anterior temporal lobes. The hippocampus transforms short-term experiences into long-term memories stored in the cortex [11], serving as a knowledge base for analogical retrieval. The left frontopolar cortex and inferior PFC aid in semantic processing and relational integration, while the right dorsolateral PFC supports inference and decision-making [56, 12]. The anterior temporal lobes further assist in linking new stimuli with prior knowledge. At the cellular level, analogical learning involves the formation of new synaptic pathways, enabling the integration and abstraction of novel concepts [13]. Inspired by these principles, we propose a brain-inspired analogical generator that emulates this mechanism for few-shot class-incremental learning.
@@ III-B1 Knowledge Base Construction
where, ${\rm{h}}_{W}(\cdot)$ is the linear classification layer and $f\left(\cdot\right)$ denotes the feature extractor. $N_{D^{0}}$ is the number of the samples of $\mathcal{D}^{0}$. $N_{C^{j}}$ denotes the number of the samples of $\mathcal{C}^{j}$
@@
5:Compute prototypes $\boldsymbol{p}^{0}$ using $f(\cdot)$triangleright Eq.(2)
@@ III-B2 Analogical Generator Training
In this stage, BiAG is trained to have the capability of generating new class weights through analogy. To achieve this goal, we take the following steps:
@@
${\rm{G}}={\varphi}_{\rm{{}_{BiAG}}}({\boldsymbol{p}^{0}_{old}},{\boldsymbol{p}^{0}_{new}},{{\rm{W}}_{old}^{0}}){\boldsymbol{p}_{{{c}_{i}}}}=\frac{1}{{{N}_{{{c}_{i}}}}}\sum\nolimits_{j=1}^{{{N}_{{{c}_{i}}}}}{f({{x}_{j,{{c}_{i}}}})}$ (3)
where ${\varphi}_{\rm{{}_{BiAG}}}(\cdot)$ denotes BiAG, and details about BiAG will be elaborated in 3.3. To ensure the stability and accuracy of $G$, we use cosine similarity between $G$ and $W_{new}^{0}$ to constrain $G$:
@@ III-B3 Weight Generation for New Classes
In this stage, both the feature extractor $f(\cdot)$ and the analogical generator BiAG ${\varphi}_{\rm{{}_{BiAG}}}(\cdot)$ are frozen to preserve the previously acquired knowledge. The goal is to generate classification weights for newly emerged classes using only a few labeled samples. This is achieved by analogically leveraging the semantic and structural knowledge encoded in the prototypes and weights of old classes.
@@
This process can be viewed as a form of cognitive analogy, where the system constructs an understanding of the new classes by relating them to known ones in the learned semantic space.
Compared to fine-tuning or storing exemplars, this strategy provides a flexible and efficient mechanism for incorporating novel categories. It avoids direct parameter updates and minimizes interference with previously learned knowledge, which is crucial for preventing catastrophic forgetting in FSCIL scenarios.
@@ III-C Brain-inspired Analogical Generator
Inspired by the prefrontal cortex’s capacity for abstract reasoning and relational mapping, as shown in Fig.3, BiAG integrates three specialized modules to emulate analogical reasoning: the Weight Self-Attention module (WSA), the Weight & Proto Analogical Attention (WPAA), and the Semantic Conversion module(SCM).
Figure 3: The architecture of the proposed Brain-Inspired Analogical Generator (BiAG). BiAG consists of stacked layers of Weight Self-Attention and Weight & Proto-Analogical Cross-Attention modules. A shared Semantic Conversion Module (SCM) transforms learnable queries between prototype and weight semantics to facilitate analogical generation of new class weights.
@@ III-C1 Semantic conversion module (SCM)
The SCM transforms the semantics of learnable queries between prototypes and weights based on Neural Collapse theory. According to Neural Collapse theory [57], during the terminal phase of deep neural network training, features of samples from the same class collapse to their class mean, and these class means (prototypes) become symmetrically arranged as a simplex ETF (Equiangular Tight Frame) centered at the origin. Simultaneously, the classification weight vectors align with their corresponding class prototypes, forming an almost identical geometric structure in the dual space—a phenomenon known as self-dual alignment. This alignment implies a tight coupling between the classification weights and the class prototypes, allowing them to be interchanged or mapped onto each other. This theoretical insight directly motivates the design of the Semantic Conversion Module (SCM) in the Brain-Inspired Analogical Generator (BiAG). By leveraging the correspondence between prototypes and classification weights revealed by Neural Collapse, SCM enables effective translation between feature space and classifier space. In essence, SCM is grounded in the understanding that class prototypes and classifier weights encode the same discriminative information, albeit in different forms.
The main function of SCM is to convert the semantics between prototypes and weights, thereby assisting the analogical learning process of BiAG. SCM employs a multi-layer perceptron (MLP) to transform the semantics of the learnable query from prototype to weight, or vice versa, to adapt to different learning stages. SCM inputs the learnable query with weight semantics after conversion into WSA for supplementary knowledge computation and inputs the learnable query with prototype semantics into WPAA for analogical learning. Additionally, SCM converts the output of the previous layer into class prototype semantics and adds it to the learnable query for updating.
III-C2 Weight Self Attention (WSA)
The Weight Self-Attention (WSA) module is designed to enhance the weights of new classes by supplementing relevant knowledge. WSA selectively focuses on important features in the input data through a self-attention mechanism, thereby improving the representational capacity of the new class weights. Additionally, WSA introduces a learnable query $\boldsymbol{q}_{L}\in\mathbb{R}^{|\mathcal{C}^{t}|\times D}$ to further enhance the efficiency of utilizing new class knowledge. The learnable query is initialized with the prototype of the new class $\boldsymbol{p}^{t}\in\mathbb{R}^{|\mathcal{C}^{t}|\times D}$ and transformed into weights by the SCM to ensure the correct semantic representation and accurate expression of new class information:
$\boldsymbol{q}_{W}={\varphi}_{\rm{{}_{SMC}}}(\boldsymbol{q}_{L})$ (7)

### v2

@@ Abstract
Few-shot class-incremental Learning (FSCIL) enables models to learn new classes from limited data while retaining performance on previously learned classes. Traditional FSCIL methods often require fine-tuning parameters with limited new class data and suffer from a separation between learning new classes and utilizing old knowledge. Inspired by the analogical learning mechanisms of the human brain, we propose a novel analogical generative method. Our approach includes the Brain-Inspired Analogical Generator (BiAG), which derives new class weights from existing classes without parameter fine-tuning during incremental stages. BiAG consists of three components: Weight Self-Attention Module (WSA), Weight & Prototype Analogical Attention Module (WPAA), and Semantic Conversion Module (SCM). SCM uses Neural Collapse theory for semantic conversion, WSA supplements new class weights, and WPAA computes analogies to generate new class weights. Experiments on miniImageNet, CUB-200, and CIFAR-100 datasets demonstrate that our method achieves higher final and average accuracy compared to SOTA methods.
publicationid: pubid: Copyright © 2025 IEEE. Personal use of this material is permitted. However, permission to use this material for any other purposes must be obtained from the IEEE by sending an email to pubs-permissions@ieee.org.
Figure 1: An example of human analogical learning mechanism and the proposed Brain-Inspired Analogical Generator (BiAG). (a) Humans can efficiently form concepts of new classes through analogical reasoning. For example, when encountering a new animal like a panda, one may relate its visual characteristics to previously known classes—such as the body shape of a bear and the black-and-white color pattern of a zebra. This analogy enables the construction of a concept of ”panda” even from limited samples. (b) Inspired by this cognitive mechanism, our BiAG module analogically generates classification weights for new classes by leveraging prototypes and weights from old classes along with a learnable query derived from new class knowledge. This allows efficient integration of new classes in few-shot incremental learning.
@@ I Introduction
Conventional deep learning models have achieved remarkable success in various computer vision and pattern recognition tasks. However, these models are typically developed under a closed-world assumption, where all classes are predefined, and ample labeled data is available for each class. In practice, this assumption rarely holds. New classes may appear continuously over time, while labeled data is often expensive or difficult to obtain. To address this limitation, Class-Incremental Learning (CIL) has been proposed to enable models to learn new classes sequentially while retaining prior knowledge. Although effective in fully supervised settings, most existing CIL methods assume sufficient data for each new class and often require storing exemplars from past sessions. These assumptions greatly limit their applicability in dynamic and data-constrained environments.
@@
In many real-world scenarios, such as identifying rare species, adapting to evolving user behavior, or updating clinical diagnosis systems, new classes often emerge under tight time constraints and with very limited data. In such settings, retraining models from scratch or storing all historical data is often infeasible. Moreover, traditional CIL methods tend to overfit to the few new samples or suffer from forgetting previously learned knowledge. To address these challenges, Few-Shot Class-Incremental Learning (FSCIL) has been proposed [1]. FSCIL requires models to learn new classes from only a handful of labeled examples while maintaining performance on old classes. Typically, the model is trained in a base session with sufficient data, followed by multiple incremental sessions with only a few examples per new class. Data from previous sessions becomes unavailable, and the model must perform classification over all seen classes. This setting emphasizes the need for methods that can effectively balance stability and plasticity under strict data and time constraints.
FSCIL faces two key challenges: overfitting due to limited new-class samples and catastrophic forgetting of old classes [1]. Existing methods fall into two classes: unified learning and sequential learning. Unified methods optimize for all classes jointly using meta-learning [2], replay [3], or dynamic networks [4] to retain prior knowledge. In contrast, sequential approaches decouple new and old class learning, often focusing on efficient representation[5][6][7]. Despite their strengths, both approaches often require fine-tuning, leading to overfitting on limited new data. Moreover, by treating new class learning and old knowledge retention as separate processes, they introduce a disconnect that hinders the balanced integration of past knowledge and new information.
@@
Inspired by the analogical learning mechanism of the human brain, we propose a novel analogical generation method that derives new class weights from existing ones. Our method sufficiently utilizes prior knowledge during incremental sessions while eliminating the need for parameter training. The framework includes a feature extractor and a Brain-Inspired Analogical Generator (BiAG). The BiAG takes the prototypes (the centric or mean feature of an identical class) of new and old classes, as well as the classification weights of old classes, as inputs to the module, and outputs the generated classification weights of the new classes. The BiAG comprises three core components: the Semantic Conversion Module (SCM), the Weight Self-Attention Module (WSA), and the Weight & Prototype Analogical Attention Module (WPAA). The SCM facilitates smooth transitions between prototypes and classification weights based on Neural Collapse theory. The WSA enhances knowledge utilization by supplementing new class weights, while the WPAA draws analogies between old and new classes, reorganizing old weights to generate accurate new class weights. Motivationally, WSA can be seen as a preparatory step that refines the new-class representation, emphasizing its salient attributes and aligning it with prior knowledge. This mirrors human cognition, where one first identifies the key features of a new concept before drawing analogies to known classes. Building upon this foundation, our framework is structured into two distinct phases: the base session training and incremental session inference phases. During the base session, the feature extractor is trained on the base classes to obtain initial classification weights and prototypes, thereby constructing the knowledge base for analogical learning. Subsequently, the BiAG is trained to cultivate the ability to generate weights for novel classes through simulated incremental learning, wherein we randomly divide the base classes into old and new classes and train the model to generate weights for new classes. In the incremental inference phase, as illustrated in Fig. 1(b), the BiAG generates new class weights by leveraging the prototypes of the new classes alongside the prototypes and weights of the old classes.
@@ II-A Class-Incremental Learning (CIL)
Class-incremental learning (CIL) enables models to continually recognize new classes while preserving knowledge of previously learned ones. The main challenge lies in catastrophic forgetting, where learning new tasks interferes with old knowledge. To address this, a variety of strategies have been developed. Regularization-based methods [13][14][15] penalize changes to important parameters to preserve past knowledge. For example, EWC [15] leverages Fisher Information to identify critical weights. More recent efforts introduce forgetting-aware directions and equilibrium constraints to improve stability [16]. Replay-based methods [17][18][19][20] retain exemplars or use generative models [21] to rehearse previous classes. To overcome memory and privacy concerns, recent approaches replace exemplars with mixed features and auxiliary classes [22] or use semantic mapping and background calibration to improve alignment [23]. Statistical sampling has also been proposed to simulate past data distributions without real samples [24]. Parameter isolation strategies [25][26][27][28] assign different parameter subsets to different tasks, avoiding interference. PackNet [26], for instance, uses weight pruning and reallocation to accommodate new tasks efficiently. Dynamic architecture methods [29][30][31][32] incrementally grow the network to handle new tasks, though often at the cost of increased complexity. Recent improvements include lightweight incremental frameworks [33] and memory-boosted transformers [34]. In addition, several works enhance representation quality to improve learning efficiency. For example, feature expansion and robust training techniques have proven effective in exemplar-free settings [35], while others leverage external semantic knowledge to guide feature learning and improve generalization [36]. CdCIL [37] proposes a curiosity-driven sample selection strategy, where uncertainty and novelty metrics are jointly used to identify informative training samples for continual adaptation.
@@ II-B Few-Shot Class-Incremental Learning (FSCIL)
Few-shot class-incremental learning (FSCIL) aims to learn new classes from few samples over time without forgetting old ones. Early works such as TOPIC [1] and CEC [4] adopt feature structure preservation or graph-based adaptation. Meta-learning-based methods like MetaFSCIL [2] and SPPR [38] improve adaptability via task sampling and prototype refinement. SPPR[38] and IDLVQ [39] enhance prototype-based learning through episodic sampling and vector quantization. To address representation degradation, FACT [6] pre-allocates virtual prototypes, while Data-free Replay [40] replays knowledge to balance class learning. LIMIT [41] and CLOM [42] employ transformers and margin constraints to align semantics and reduce overfitting. GKEAL [43] uses Gaussian kernels for analytic learning, and BiDist [3] distills knowledge from dual teachers. SVAC [7] enhances generalization via virtual classes. Dynamic model-based approaches include SoftNet [44], WaRP [45], and KRRM [46], each tailoring model structures for incremental adaptation. NC-FSCIL [47] pre-assigns classifiers based on Neural Collapse, while ALICE [48] and OrCo [49] improve intra-class compactness and generalization via angular loss or orthogonality. Comp-FSCIL [50] further exploits compositional knowledge to enhance few-shot adaptation. YourSelf [51] employs a lightweight Feature Rectification module to refine final features using intermediate ones, guided by relation transfer losses and multi-layer ensemble. Recent methods like PCL [52] and LRT [53] use pretrained CLIP models for feature extraction, but risk leaking novel class information via shared embeddings. While ALFSCIL [54] explores analogical learning, it only leverages prototypes and still relies on fine-tuning. In contrast, our method, BiAG, fully utilizes both prototypes and old class weights to analogically generate new classifiers in a scalable and tuning-free manner.
@@ II-C Analogical Learning Mechanism
Analogical learning is a fundamental aspect of human cognition, allowing individuals to transfer knowledge from familiar domains to novel situations by recognizing relational similarities. This mechanism supports efficient and flexible learning and has been identified by cognitive science as a key driver of human intelligence [8, 9, 55]. According to Gentner’s Structure-Mapping Theory [56], analogical reasoning involves retrieving relevant experiences, aligning relational structures (mapping), projecting inferences, evaluating consistency, and re-representing knowledge when necessary.
Neurologically, analogical reasoning is supported by a distributed network involving the hippocampus, prefrontal cortex (PFC), and anterior temporal lobes. The hippocampus transforms short-term experiences into long-term memories stored in the cortex [10], serving as a knowledge base for analogical retrieval. The left frontopolar cortex and inferior PFC aid in semantic processing and relational integration, while the right dorsolateral PFC supports inference and decision-making [11]. At the cellular level, learning relies on structural synaptic plasticity, where the formation and stabilization of new synaptic connections and dendritic spine modifications contribute to the consolidation of long-term memory [12]. Inspired by these principles, we propose a brain-inspired analogical generator that emulates this mechanism for few-shot class-incremental learning.
@@ III-B1 Knowledge Base Construction
where, ${\rm{h}}_{W}(\cdot)$ is the linear classification layer and $f\left(\cdot\right)$ denotes the feature extractor. $N_{D^{0}}$ is the number of the samples of $\mathcal{D}^{0}$.
@@
5:Compute prototypes $\boldsymbol{p}^{0}$ using $f(\cdot)$ $\triangleright$ Eq.(2)
@@ III-B2 Analogical Generator Training
In this stage, BiAG is trained to have the capability of generating new class weights through analogy. To achieve this goal, steps are as follows:
@@
${\rm{G}}={\varphi}_{\rm{{}_{BiAG}}}({\boldsymbol{p}^{0}_{old}},{\boldsymbol{p}^{0}_{new}},{{\rm{W}}_{old}^{0}})$ (3)
where ${\varphi}_{\rm{{}_{BiAG}}}(\cdot)$ denotes BiAG. To ensure the stability and accuracy of $G$, we use cosine similarity between $G$ and $W_{new}^{0}$ to constrain $G$:
@@ III-B3 Weight Generation for New Classes
In this stage, both the feature extractor $f(\cdot)$ and the analogical generator BiAG ${\varphi}_{\rm{{}_{BiAG}}}(\cdot)$ are frozen to preserve the previously acquired knowledge. The goal is to generate classification weights for new classes using only a few labeled samples. This is achieved by analogically leveraging the semantic and structural knowledge encoded in the prototypes and weights of old classes.
@@
This process can be viewed as a form of cognitive analogy, where the system understands new classes by linking them to known ones in the learned semantic space.
Compared to fine-tuning or storing exemplars, this strategy provides a flexible and efficient mechanism for incorporating novel classes. It avoids direct parameter updates and minimizes interference with previously learned knowledge, which is crucial for preventing catastrophic forgetting in FSCIL scenarios.
@@ III-C Brain-Inspired Analogical Generator
Inspired by the prefrontal cortex’s capacity for abstract reasoning and relational mapping, as shown in Fig.3, BiAG integrates three specialized modules to emulate analogical reasoning: the Weight Self-Attention module (WSA), the Weight & Prototype Analogical Attention (WPAA), and the Semantic Conversion module (SCM).
Figure 3: The architecture of the proposed Brain-Inspired Analogical Generator (BiAG). BiAG consists of stacked layers of Weight Self-Attention and Weight & Prototype Analogical Cross-Attention modules. A shared Semantic Conversion Module (SCM) transforms learnable queries between prototype and weight semantics to facilitate analogical generation of new class weights.
@@ III-C1 Semantic Conversion Module (SCM)
The SCM transforms learnable queries between prototype semantics and weight semantics based on the Neural Collapse (NC) theory [57]. According to NC, during the terminal phase of network training, features and classifier weights exhibit the following properties:
  • NC1 (Feature Collapse): Let $\mu_{k,i}\in\mathbb{R}^{d}$ denote the feature of the $i$-th sample in class $k$, and let $n_{k}$ be the number of samples in class $k$. The class mean (prototype) is defined as $\mu_{k}=\frac{1}{n_{k}}\sum_{i}\mu_{k,i}$, and all features of class $k$ collapse to $\mu_{k}$.
  • NC2 (Simplex ETF of Class Means): Let $\mu_{G}=\frac{1}{K}\sum_{k=1}^{K}\mu_{k}$ denote the global mean over $K$ classes. The centered class means $\widetilde{\mu}_{k}=\mu_{k}-\mu_{G}$ form a Simplex Equiangular Tight Frame (ETF), satisfying $\langle\widetilde{\mu}_{k},\widetilde{\mu}_{k^{\prime}}\rangle=-\frac{c^{2}}{K-1}$ for $k\neq k^{\prime}$, with $\|\widetilde{\mu}_{k}\|=c$.
  • NC3 (Self-Dual Alignment): Let $w_{k}\in\mathbb{R}^{d}$ denote the classifier weight for class $k$. The weight directions align with the normalized prototypes, $\frac{w_{k}}{\|w_{k}\|}=\frac{\widetilde{\mu}_{k}}{\|\widetilde{\mu}_{k}\|}$.
  • NC4 (Nearest-Mean Prediction): A sample is classified by the nearest class mean, $\hat{y}=\arg\max_{k}\langle w_{k},\mu\rangle=\arg\min_{k}\|\mu-\mu_{k}\|^{2}$.
These properties indicate that class prototypes $\{\mu_{k}\}$ and classifier weights $\{w_{k}\}$ share the same Simplex-ETF geometry. Therefore, there exists an affine mapping between prototypes and weights:
$w_{k}=s\,{\bf R}\,\widetilde{\mu}_{k}={\bf A}\mu_{k}+{\bf b},$ (7)
where $s>0$ is a scalar, ${\bf R}\in\mathbb{R}^{d\times d}$ is an orthogonal matrix, ${\bf A}=s{\bf R}$, and ${\bf b}=-s{\bf R}\mu_{G}$. This affine relation preserves the NC2–NC3 Simplex-ETF geometry up to a common rotation and scale. The SCM is implemented as a multi-layer perceptron (MLP) that can exactly represent this affine mapping, thereby preserving the NC2–NC3 geometry between prototypes and weights. Within BiAG, SCM fulfills two key roles: it first converts the learnable query into weight semantics for the WSA module to generate supplementary knowledge, and then converts the intermediate outputs into prototype semantics for the WPAA, ensuring semantic consistency and enabling effective analogical weight generation.
III-C2 Weight Self-Attention (WSA)
The WSA module is designed to enhance the weights of new classes by supplementing relevant knowledge. WSA selectively focuses on important features in the input data through a self-attention mechanism, thereby improving the representational capacity of the new class weights. Additionally, WSA introduces a learnable query $\boldsymbol{q}_{L}\in\mathbb{R}^{|\mathcal{C}^{t}|\times D}$ to further enhance the efficiency of utilizing new class knowledge. The learnable query is initialized with the prototype of the new class $\boldsymbol{p}^{t}\in\mathbb{R}^{|\mathcal{C}^{t}|\times D}$ and transformed into weights by the SCM to ensure the correct semantic representation and accurate expression of new class information:
$\boldsymbol{q}_{W}={\varphi}_{\rm{{}_{{SCM}}}}(\boldsymbol{q}_{L})$ (8)